How do you find the area of a Pentagon with the side lenths of 2?
Answer by jedidiah(6)

area of triangle (1/5 area of polygon) = (1/2)bh
b=2
angle of each vertex:
(5-2)(180 degrees)/5 vertices = 108 degrees
angle of triangle opposite h:
108/2 = 54 degrees
height of triangle:
h= 1*tan(54 degrees)= 1.38
area of triangle= (1/2)*2*1.38= 1.38
area of pentagon:
5*1.38= 6.9 units^2